6b) The following table gives the hours spent by 50 students in the playground in a week:

Hours 0 2 4 6 8 10

No. of students 10 6 8 6 10 10

Calculate the mean median and standard deviation of the hours required

by students.


1
Expert's answer
2021-05-07T13:38:02-0400

Table of this question can be:




Mean "= \dfrac{232}{50} = 4.64"


Standard Deviation "= \sqrt{\dfrac{\sum(x_i-\bar x)^2}{N}} = \sqrt{\dfrac{70.73}{50}} = 1.18"


Median "= \dfrac{25^{th}+26^{th}}{2} = \dfrac{6+6}{2} = 6"